OpenStudy (anonymous):

Simplify the terms. (2.5 x 10^-3) x (3 x 10^-4) (9.6 x 10^26) divided by (3.2 x 10^9) (2.5 x 10^8) x (7 x 10^5) (36.16 x 10^-7) divided by (4 x 10^-20)

OpenStudy (anonymous):

all of these are 10 to the power of something, that makes it easy

OpenStudy (anonymous):

for example\( \Large \frac{(2.8*10^{12})}{(2*10^3)}=\frac{2.8}{2}*10^{12-3}=1.4*10^9\)

OpenStudy (anonymous):

using that example try solving the other 4